OpenStudy (anonymous):

write an equation in the form Ax+By=C of the line that passes through the point (-2,3) and is parallel to 7x-2y=14 how do i do this?

OpenStudy (anonymous):

okay first take 7x-2y+14 and put it in y=mx+b form. do you know how to do that?

OpenStudy (anonymous):

sorry =14

OpenStudy (anonymous):

okayi did that and i got y= 7/2x-7

OpenStudy (anonymous):

okay good..now now plug in the coordinated in y=7/2x-b x=-2 y=3

OpenStudy (anonymous):

coordinates*

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

y=-7x?

OpenStudy (anonymous):

actually when you plug in..you are trying to find b. so what is b?

OpenStudy (anonymous):

oh wait -7x+2y=20?

OpenStudy (anonymous):